OpenStudy (anonymous):

what is the percent of 480 out if 1200 what is the percent of 480 out if 1200 @Mathematics

OpenStudy (anonymous):

can some one please work this problem for me what is the percent of 480 out of 1200

OpenStudy (amistre64):

percents are just part divided by total; and then you multiply by 100 to get the proper format

OpenStudy (amistre64):

480/1200 48/120 4/10 = .4 .4 * 100 = 40, so 40%

hero (hero):

2/5 = 40%